OpenStudy (anonymous):

Mr. B needs to make 10L of 42% solution. The solutions are available at 30% of solution x and 50% of solution y. how many litres of each solution must be mixed to make the 42% solution? (I need 2 equations. don't solve, just 2 equations)

OpenStudy (mertsj):

x+y=10 .30x+.50y=.42(10)

OpenStudy (anonymous):

that was the same answer, i'm just wondering why isn't is .3x+.5y=.42? why is it 4.2?

OpenStudy (whpalmer4):

in the second equation, you're computing the amounts of the actual "stuff" in the solution. x liters of 30% solution gives you 0.3x of the solute. etc.

OpenStudy (whpalmer4):

and you're trying to make a 42% solution, 10 liters, so there will be 0.42*10 = 4.2 l of "stuff" in the final mix

OpenStudy (anonymous):

oh i get it, you always have to do x10 to the total amount?

OpenStudy (mertsj):

Because the second equation is based on the amount of solute in each solution. x liters of 30% solution contains .3x l of solute. 10 liters of 42% solution contains .42(10) liters of solute

OpenStudy (whpalmer4):

that's only because the problem asks to make 10 liters. if it asked for 23 liters, the second equation would be 0.30x + 0.50y = 0.42(23)

OpenStudy (whpalmer4):

(and the first would be x + y = 23)

OpenStudy (anonymous):

ohhhhh thank you!